Tyrion:

Show all work to identify the asymptotes and zero of the function f of x equals 3 x over quantity x squared minus 9. I HAVE BEEN TRYNA DO THIS PROBLEM BY MYSELF FOR 3 DAYS NOW KAN SUMONE PLZ HELP

AZ:

\( f(x) = \dfrac{3x}{x^2 - 9}\) is this it?

Tyrion:

yes

AZ:

The vertical asymptotes can be found when you set the denominator = 0 do you know how to factor \( x^2 - 9\) hint: \( a^2 - b^2 = (a+b)(a-b)\)

Tyrion:

no i dont know nun of dis stuff,,she said if we kan answer this question she will give us bonus points,,but we never did this math

AZ:

But can you factor \( x^2 - 9\) hint: \( 9 = 3^2\)

Tyrion:

ion rlly kare about the points but i juss wanna see how to do it for future problems

AZ:

So tell her that to find the vertical asymptote, you have to set the denominator equal to 0 and once you factor x^2 - 9, you'll have your answer in no time

AZ:

and then the bonus points shall be yours, my friende

Tyrion:

(x-3) (x+3)

AZ:

Perfect!! so remember we said we're setting the denominator equal to 0 (x^2 - 9) = 0 (x+3)(x-3) = 0 so now we have x + 3 = 0 and x - 3 = 0 what two numbers do you get when you solve for x?

Tyrion:

-3 and 3???

snowflake0531:

Yes~ Which means that x cannot equal -3 or 3, so they are your asymptotes

AZ:

Exactly! So those are your asymptotes x = 3 and x = -3 keep the x equals to part

AZ:

Now we have one last thing- we need to find the zeroes of the function

Tyrion:

wat is dat

snowflake0531:

Zeroes are like the x-intercepts

AZ:

the zero of the function is the x-intercept which is when y = 0 so take a look at your function and put y = 0 \( 0 = \dfrac{3x}{x^2 -9}\) to make it equal to 0, we need the NUMERATOR to equal 0 if the denominator equals 0 then it would be undefined so what times 3 will give you 0?

Tyrion:

0

AZ:

So that's the zero of your function! On the graph, it would be when it crosses the x-axis at (0, 0)

AZ:

you can see the asymptotes as x = -3 and x = 3 https://www.desmos.com/calculator/ikl8cvsnq6

Tyrion:

thanx

AZ:

You're welcome!